# Hug Reports: Supporting Expression of Appreciation between Users and Contributors of Open Source Software Packages

### v1

@@ 5.4.2. Appreciation as an act of expressive communication (RQ2-U, RQ2-C)
While some personal notes did mention these aspects (5.1.1), many users struggled to find something specific to say, in the moment, which led to them writing generic notes of appreciation or prevented them from writing personal notes altogether. U7 mentioned he was uncertain about what would make “good communication in this relationship between the user and contibutor”. Other users mentioned how it was a “bit tricky to explain why, I’m actually grateful” (U12) and how they sometimes “didn’t have anything specific to say” (U6). U9 suggested: “it was difficult to know what to say other than thanks. It’s a lot easier to write something when things go wrong.”
When considering thanks as a personal expression of gratitude, some users described how their experiences were affected by the size of the anticipated audience of their thanks (U6, U12). In the context of larger packages, these users felt that thanks scoped to the entire package were less meaningful than thanks scoped to individual modules or function.
5.5. Contributors’ reactions were influenced by how much they felt they had contributed to the object that was thanked (RQ2-C)
Contributors felt that thanks was undeserved if felt they had not substantially contributed to the thanked code. Even if contributors had made code edits to the part of the package that was thanked, they felt it was undeserved if they did not feel a sense of ownership. For instance, in their survey response C20 described:
“The thanks I received were for a Python package that I happily used myself and contributed a minimum amount to (two lines of code around two years ago or such). Therefore, I don’t believe to deserve praise for my contributions. However, delivering and receiving thanks gives me a good feeling, as I am happy for the actual contributors to be acknowledged.” (C20)
@@
“It depends on my [sense of] ownership of that project. If that is above a certain threshold, then that thanks is directly meaningful to me. So like something like like 5% or 10% would be that threshold that comes to my mind that that if I’ve written more than 10% of the code in the project, I would want to know when when someone is thanking the project.”(C21)
This is not necessarily a critical failure of Hug Reports. Some contributors felt that these “misfires” were not an issue: “spreading the thanks signal out into the world is an example of something that, even if it misfires, it doesn’t hurt and so, I don’t think that the misfires are necessarily bad” (C13). C18 described how she “wouldn’t feel weird” if thanks were misdirected at her, further adding: “it would be really easy for me to redirect it to the right person”. Still, we recognize that better designed heuristics for identifying relevant contributors can enhance the meaningfulness of the thanks.
@@ 5.6.2. Opportune moments
Reflecting on patterns in when they tend to express appreciation, users pointed out ways in which the system could harness opportune moments. U1 suggested: “every time you open or close a file maybe you could get a pop up which says that, ‘oh, in this file you use these things”’. U7 had a similar suggestion, mentioning how he would like such notifications to emphasize new features he used that week. U9 and U10 mentioned how they would like to view statistics of their use, at times that they could control (e.g. immediately after a specific recurring meeting), either via a notification or dashboard. C12 suggested that the end of project cycles, if users were themselves publishing work to GitHub, it could be a opportune moment to remind users of the packages they were using: “GitHub, has a feature where you say like, make a release right? You could tie an action into that where it would say, ‘Oh, you’re making a release. Here are all the [packages] you used in your project in this release”’ (C12).
@@ 5.6.3. Form and frequency of notifications
Contributors described how the form in which Hug Reports were aggregated and presented could be broadened to support different use cases: “the psychological benefits of it would be garnered more by sending the emails out, and the economic benefits might be garnered better by putting it up on a website and sharing the link around” (C13). Decision-making and fund raising activities could benefit from more aggregated statistics: “from the perspective of the project, we wouldn’t really care about the personal notes, because we would aggregate over them anyway” (C8). Contributors also suggested venues to make the aggregate data public, such as in project “release notes” (C8) and in individuals’ “GitHub profiles” (C27). C20 and C18 also felt having a way to display on the project’s GitHub page, for example as “a GitHub badge” (C18) would be nice: “on GitHub, I look at the number of give stars, the number of different issues and so, it seems the number of happy users, would also be a good metric” (C18)
@@ 5.6.4. Scaffolding for personal effort
Both users and contributors suggested it would be useful to have prompts that could help users write more specific personal notes: “maybe giving some instruction of like to users like, here’s an example. What you can write like here was my used case, and why it was useful.” (U9). C19 suggested having something similar to product reviews that ask users “what’s good about it? Is the make good? Is the design good?”. U3 suggested it could be useful to be able to “send pictures, you know, especially if things are hardware related”. In his survey response, C26 suggested having discrete pre-determined categories: “maybe a selection from a few pre-determined thank-types. e.g ‘I am using it for everything, thanks!’ or ‘Saved me a ton of time in my current project”’.
@@ 6. Discussion
Through our deployment, we found that by lowering the procedural effort and by reminding users of the contributors, Hug Reports encouraged appreciation in ways that were meaningful to users and contributors. Other takeaways included patterns in when users expressed appreciation, how appreciation took on two meanings, and how contributors’ reactions to the appreciation were influenced by their perceived level of impact. In addition to this, users and contributors proposed several ideas for how appreciation could be better supported. Here, we begin by discussing how our work contributes new knowledge on designing appreciation systems in peer production by exploring the implications of re-orienting appreciation around abstractions of use. Then, we discuss how our work extends prior literature on appreciation in open source by contributing insights into how appreciation is experienced and expressed by users in practice. Finally, we discuss limitation of our work and directions for future work.
@@ 6.1. Opportunity and limitation of re-orienting appreciation around abstractions of use
Our approach tried to enable users to express appreciation in terms of the abstractions they are exposed to—the package or its modules—rather than lower-level units of contribution such as individual commits or pull requests. This made it unique from Wikipedia’s “Thanks” system that captures appreciation towards individual edits. Our findings show how re-orienting appreciation around packages and modules meant it aligned more with how users experience appreciation and made it easier for them to express it. At the same time, shifting appreciation away from low-level units of contribution limited the extent to which appreciation could serve as a form of individual recognition. Here, we discuss the opportunity and limitation of this approach.
@@ 6.3.1. Limitations in the kinds of contributions considered
Our inquiry was limited to scenarios where open source software use occurred through packages. Future work can explore whether and how these approaches can be extended to other kinds of open source software such as end-user applications and cookie cutter templates. Further, we chose to focus on Python and JavaScript packages because they are commonly used languages with large package ecosystems, and are also languages we are most familiar with. Future work can explore extending this approach to other programming languages/ecosystems. By relying on code activity traces, our work supported appreciation of contributors who made code contributions to projects. However, this approach overlooks many important non-code contributions such as documentation, governance, fund-raising, community-building. Hence, there is an opportunity for future research to investigate how appreciation can be extended to non-code contributions.
@@ 7. Conclusion
Contributors of open source software packages rarely receive appreciation from users. In this paper, we observed how appreciation can be limited by the fact that where users might feel appreciation (in their development environment) and what they might feel appreciation towards (a package, its modules, or its functions) is detached from where contribution activities occur (GitHub) and what its units are (individual commits or pull requests). We described a field study of the Hug Reports technology probe that provided users with a communication affordance within the code editor, and allowed them to express appreciation in terms of the abstractions they are exposed to (package, modules). Our findings showed how Hug Reports encouraged appreciation in ways that were meaningful to users and contributors, how appreciation was interpreted both as a measure of utility and as an act of expressive communication, and that contributors’ reactions to appreciation were influenced by how much they felt they had contributed to what was thanked. In addition to this, our study revealed patterns in when users expressed appreciation. We synthesized these findings into implications for developing appreciation systems in open source in particular, and peer production communities more generally.

### v2

@@ 5.4.2. Appreciation as an act of expressive communication
While some personal notes did mention these aspects (5.1.1), many users struggled to find something specific to say, in the moment, which led to them writing generic notes of appreciation or prevented them from writing personal notes altogether. U7 mentioned he was uncertain about what would make “good communication in this relationship between the user and contributor”. Other users mentioned how it was a “bit tricky to explain why, I’m actually grateful” (U12) and how they sometimes “didn’t have anything specific to say” (U6). U9 suggested: “it was difficult to know what to say other than thanks. It’s a lot easier to write something when things go wrong.”
When considering thanks as a personal expression of gratitude, some users described how their experiences were affected by the size of the anticipated audience of their thanks (U6, U12). In the context of larger packages, these users felt that thanks scoped to the entire package were less meaningful than thanks scoped to individual modules or functions.
5.5. Contributors’ reactions were influenced by how much they felt they had contributed to the object that was thanked (RQ2)
Contributors felt that thanks was undeserved if felt they had not substantially contributed to the thanked code. Even if contributors had made code edits to the part of the package that was thanked, they felt it was undeserved if they did not feel a sense of ownership. For instance, in their survey response, C20 described:
“The thanks I received were for a Python package that I happily used myself and contributed a minimum amount to (two lines of code around two years ago or such). Therefore, I don’t believe I deserve praise for my contributions. However, delivering and receiving thanks gives me a good feeling, as I am happy for the actual contributors to be acknowledged.” (C20)
@@
“It depends on my [sense of] ownership of that project. If that is above a certain threshold, then that thanks is directly meaningful to me. So something like 5% or 10% would be that threshold that comes to my mind. If I’ve written more than 10% of the code in the project, I would want to know when someone is thanking the project.”(C21)
This is not necessarily a critical failure of Hug Reports. Some contributors felt that these “misfires” were not an issue: “spreading the thanks signal out into the world is an example of something that, even if it misfires, it doesn’t hurt and so, I don’t think that the misfires are necessarily bad” (C13). C18 described how she “wouldn’t feel weird” if thanks were misdirected at her, further adding: “it would be really easy for me to redirect it to the right person”. Still, we recognize that better-designed heuristics for identifying relevant contributors can enhance the meaningfulness of the thanks.
@@ 5.6.2. Opportune moments
Reflecting on patterns in when they tend to express appreciation, users pointed out ways in which the system could harness opportune moments. U1 suggested: “every time you open or close a file maybe you could get a pop up which says that, ‘oh, in this file you use these things”’. U7 had a similar suggestion, mentioning how he would like such notifications to emphasize new features he used that week. U9 and U10 mentioned how they would like to view statistics of their use, at times that they could control (e.g. immediately after a specific recurring meeting), either via a notification or dashboard. C12 suggested that if users were themselves publishing work to GitHub, the end of project cycles could be opportune moments to remind users of the packages they were using: “GitHub, has a feature where you say like, make a release right? You could tie an action into that where it would say, ‘Oh, you’re making a release. Here are all the [packages] you used in your project in this release”’ (C12).
@@ 5.6.3. Form and frequency of notifications
Contributors described how the form in which Hug Reports were aggregated and presented could be broadened to support different use cases: “the psychological benefits of it would be garnered more by sending the emails out, and the economic benefits might be garnered better by putting it up on a website and sharing the link around” (C13). Decision-making and fund-raising activities could benefit from more aggregated statistics: “from the perspective of the project, we wouldn’t really care about the personal notes, because we would aggregate over them anyway” (C8). Contributors also suggested venues to make the aggregate data public, such as in project “release notes” (C8) and in individuals’ “GitHub profiles” (C27). C20 and C18 also felt having a way to display on the project’s GitHub page, for example as “a GitHub badge” (C18) would be nice: “on GitHub, I look at the number of stars, the number of different issues and so, it seems the number of happy users, would also be a good metric” (C18)
@@ 5.6.4. Scaffolding for personal effort
Both users and contributors suggested it would be useful to have prompts that could help users write more specific personal notes: “maybe giving some instruction of like to users like, here’s an example. What you can write like here was my used case, and why it was useful.” (U9). C19 suggested having something similar to product reviews that ask users “what’s good about it? Is the make good? Is the design good?”. U3 suggested it could be useful to be able to “send pictures, you know, especially if things are hardware related”. In his survey response, C26 suggested having discrete pre-determined categories: “maybe a selection from a few pre-determined thank-types. e.g. ‘I am using it for everything, thanks!’ or ‘Saved me a ton of time in my current project”’.
@@ 5.6.5. Heuristics for notifying contributors
5.7. Summary of findings
Addressing RQ1, our deployment found that by lowering the procedural effort and by reminding users of the contributors, Hug Reports encouraged appreciation in ways that were meaningful to users and contributors (5.2, 5.1.2). Addressing RQ2, we described trends in usage (5.1.1), and contributor’s reactions (5.1.2). Further addressing RQ2, the interviews revealed patterns in when users expressed appreciation (5.3), how appreciation took on two meanings (5.4), and how contributors’ reactions were influenced by their perceived level of involvement (5.5). Finally, addressing RQ3, we presented ideas proposed by users and contributors for how appreciation could be better supported (5.6).
@@ 6. Discussion
Here, we begin by discussing how our work contributes new knowledge on designing appreciation systems in peer production by exploring the implications of re-orienting appreciation around abstractions of use. Then, we discuss how our work extends literature on appreciation in open source by uncovering how appreciation is experienced and expressed by users in practice. Finally, we discuss limitations of our work and directions for future work.
@@ 6.1. Opportunity and limitation of re-orienting appreciation around abstractions of use
Our approach tried to enable users to express appreciation in terms of the abstractions they are exposed to—the package or its modules—rather than lower-level units of contribution such as individual commits or pull requests. This made it unique from Wikipedia’s “Thanks” system which captures appreciation towards individual edits. Our findings show how re-orienting appreciation around packages and modules meant it aligned more with how users experience appreciation and made it easier for them to express it. At the same time, shifting appreciation away from low-level units of contribution limited the extent to which appreciation could serve as a form of individual recognition. Here, we discuss the opportunity and limitations of this approach.
@@ 6.3.1. Limitations in the kinds of contributions considered
Our inquiry was limited to scenarios where open source software use occurred through packages. Future work can explore whether and how these approaches can be extended to other kinds of open source software such as end-user applications and cookie-cutter templates. Further, we chose to focus on Python and JavaScript packages because they are commonly used languages with large package ecosystems, and are also languages we are most familiar with. Future work can explore extending this approach to other programming languages/ecosystems. By relying on code activity traces, our work supported appreciation of contributors who made code contributions to projects. However, this approach overlooks many important non-code contributions such as documentation, governance, fund-raising, and community-building. Hence, there is an opportunity for future research to investigate how appreciation can be extended to non-code contributions.
@@ 7. Conclusion
Contributors of open source software packages rarely receive appreciation from users. In this paper, we observed how appreciation can be limited by the fact that where users might feel appreciation (in their development environment) and what they might feel appreciation towards (a package, its modules, or its functions) is detached from where contribution activities occur (GitHub) and what its units are (individual commits or pull requests). We described a field study of the Hug Reports technology probe that provided users with a communication affordance within the code editor and allowed them to express appreciation in terms of the abstractions they are exposed to (package, modules). Our findings showed how Hug Reports encouraged appreciation in ways that were meaningful to users and contributors, how appreciation was interpreted both as a measure of utility and as an act of expressive communication, and that contributors’ reactions to appreciation were influenced by how much they felt they had contributed to what was thanked. In addition to this, our study revealed patterns in when users expressed appreciation. We synthesized these findings into implications for developing appreciation systems in open source in particular, and peer production communities more generally.